2060 2061 207C 231E 231F 230F

α \alpha  θ \theta  o o  τ \tau
β \beta  ϑ \vartheta π \pi  υ \upsilon
γ \gamma  γ \gamma  ϖ \varpi  ϕ \phi
δ \delta  κ \kappa  ρ \rho  φ \varphi
ϵ \epsilon  λ \lambda  ϱ \varrho  χ \chi
ε \varepsilon μ \mu  σ \sigma  ψ \psi
ζ \zeta  ν \nu  ς \varsigma ω \omega
η \eta  ξ \xi
Γ \Gamma  Λ \Lambda  Σ \Sigma  Ψ \Psi
Δ \Delta  Ξ \Xi  Υ \Upsilon Ω \Omega
Θ \Theta  Π \Pi  Φ \Phi

Table 1: Greek Letters

± \pm  ∩ \cap  ⋄ \diamond  ⊕ \oplus
∓ \mp  ∪ \cup  △ \bigtriangleup  ⊖ \ominus
× \times ⊎ \uplus  ▽ \bigtriangledown ⊗ \otimes
÷ \div  ⊓ \sqcap  ◁ \triangleleft  ⊘ \oslash
∗ \ast  ⊔ \sqcup  ▷ \triangleright  ⊙ \odot
⋆ \star  ∨ \vee  ⊲ \lhd^{b}  ◯ \bigcirc
∘ \circ  ∧ \wedge  ⊳ \rhd^{b}  † \dagger
• \bullet \ \setminus ⊴ \unlhd^{b}  ‡ \ddagger
· \cdot  ≀ \wr  ⊵ \unrhd^{b}  ⨿ \amalg
+ +  − -

b Not predefined in a format based on basefont.tex. Use one of the style options

           oldlfont, newlfont, amsfonts or amssymb.

Table 2: Binary Operation Symbols

≤ \leq  ≥ \geq  ≡ \equiv ⊧ \models
≺ \prec  ≻ \succ  ∼ \sim  ⊥ \perp
⪯ \preceq  ⪰ \succeq  ≃ \simeq | \mid
≪ \ll  ≫ \gg  ≍ \asymp ∥ \parallel
⊂ \subset  ⊃ \supset  ≈ \approx ▷◁ \bowtie
⊆ \subseteq  ⊇ \supseteq    ≅ \cong  ⋈ \Join^{b}
⊏ \sqsubset^{b} ⊐ \sqsupset^{b} ≠ \neq  ⌣ \smile
⊑ \sqsubseteq ⊒ \sqsupseteq  ≐ \doteq ⌢ \frown
∈ \in  ∋ \ni  ∝ \propto = =
⊢ \vdash  ⊣ \dashv  < <  > >
: :

b Not predefined in a format based on basefont.tex. Use one of the style options

           oldlfont, newlfont, amsfonts or amssymb.

Table 3: Relation Symbols

   , ,  ; ;  : \colon . \ldotp · \cdotp

Table 4: Punctuation Symbols

← \leftarrow  ⟵ \longleftarrow  ↑ \uparrow
⇐ \Leftarrow  ⟸ \Longleftarrow  ⇑ \Uparrow
→ \rightarrow  ⟶ \longrightarrow  ↓ \downarrow
⇒ \Rightarrow  ⟹ \Longrightarrow  ⇓ \Downarrow
↔ \leftrightarrow  ⟷ \longleftrightarrow ↕ \updownarrow
⇔ \Leftrightarrow  ⟺ \Longleftrightarrow ⇕ \Updownarrow
↦ \mapsto  ⟼ \longmapsto  ↗ \nearrow
↩ \hookleftarrow  ↪ \hookrightarrow  ↘ \searrow
↼ \leftharpoonup  ⇀  \rightharpoonup  ↙ \swarrow
↽ \leftharpoondown  ⇁  \rightharpoondown  ↖ \nwarrow
⇀ ↽ \rightleftharpoons ⇝  \leadsto^{b}

b Not predefined in a format based on basefont.tex. Use one of the style options

           oldlfont, newlfont, amsfonts or amssymb.

Table 5: Arrow Symbols 1

. . . \ldots · · · \cdots     ⋮ \vdots  ... \ddots
ℵ \aleph ′  \prime  ∀ \forall  ∞ \infty
h̄ \hbar ∅ \emptyset ∃ \exists  □ \Box^{b}
ı  \imath ∇ \nabla  ¬ \neg  ◇ \Diamond^{b}
ȷ \jmath √ \surd  ♭ \flat  △ \triangle
ℓ \ell  ⊤ \top  ♮ \natural  ♣ \clubsuit
℘ \wp  ⊥ \bot  ♯ \sharp  ♢ \diamondsuit
ℜ \Re  ∥ \|  \ \backslash ♡ \heartsuit
ℑ \Im     ^{}\angle  ∂ \partial  ♠ \spadesuit
℧ \mho^{b} .  .  | |

b Not predefined in a format based on basefont.tex. Use one of the style options

            oldlfont, newlfont, amsfonts or amssymb.

Table 6: Miscellaneous Symbols

         ∑ \sum  ⋂ \bigcap  ⨀ \bigodot
         ∏ \prod  ⋃ \bigcup  ⨂ \bigotimes
         ∐ \coprod ⨆ \bigsqcup ⨁ \bigoplus
         ∫ \int  ⋁ \bigvee  ⨄ \biguplus
         ∮ \oint  ⋀ \bigwedge

Table 7: Variable-sized Symbols

 \arccos  \cos   \csc  \exp  \ker     \limsup  \min  \sinh
 \arcsin  \cosh  \deg  \gcd  \lg      \ln      \Pr   \sup
 \arctan  \cot   \det  \hom  \lim     \log     \sec  \tan
 \arg     \coth  \dim  \inf  \liminf  \max     \sin  \tanh

Table 8: Log-like Symbols

 ( (  ) )  ↑ \uparrow  ⇑ \Uparrow
 [ [  ] ]  ↓ \downarrow  ⇓ \Downarrow
 { \{  } \}  ↕ \updownarrow ⇕ \Updownarrow
 ⌊ \lfloor ⌋ \rfloor ⌈ \lceil  ⌉ \rceil
 ⟨ \langle ⟩ \rangle / /  \ \backslash
 | |  ∥ \|

Table 9: Delimiters

} \rmoustache  { \lmoustache  } \rgroup     { \lgroup
| \arrowvert   ‖ \Arrowvert   | \bracevert

Table 10: Large Delimiters

â \hat{a}    á \acute{a}  ā \bar{a}  ȧ \dot{a}   ă \breve{a}
ǎ \check{a}  à \grave{a}  a⃗ \vec{a}  ä \ddot{a}  ã \tilde{a}

Table 11: Math mode accents

             ãbc  \widetilde{abc}      âbc  \widehat{abc}
             \overleftarrow{abc}  \overleftarrow\{abc\}  \overrightarrow{abc}  \overrightarrow\{abc\}
             \overline{abc}  \overline\{abc\}       \underline{abc}  \underline\{abc\}
             \overbrace{abc} \overbrace\{abc\}      \underbrace{abc} \underbrace\{abc\}
             \sqrt{abc} \sqrt\{abc\}            \sqrt[n]{abc} \sqrt[n]\{abc\}
             f′  f'                    \frac{abc}{xyz}  \frac\{abc\}\{xyz\}

Table 12: Some other constructions

                 \ulcorner     \urcorner     \llcorner     \lrcorner

Table 13: AMS Delimiters

\dashrightarrow          \dashleftarrow         \leftleftarrows       \leftrightarrows
\Lleftarrow              \twoheadleftarrow      \leftarrowtail        \looparrowleft
\leftrightharpoons       \curvearrowleft        \circlearrowleft      \Lsh
\upuparrows              \upharpoonleft         \downharpoonleft      \multimap
\leftrightsquigarrow     \rightrightarrows      \rightleftarrows      \rightrightarrows
\rightleftarrows         \twoheadrightarrow     \rightarrowtail       \looparrowright
⇀ ↽ \rightleftharpoons       \curvearrowright       \circlearrowright     \Rsh
\downdownarrows          \upharpoonright        \downharpoonright     \rightsquigarrow

Table 14: AMS Arrows

        \nleftarrow          \nrightarrow         \nLeftarrow     \nRightarrow
        \nleftrightarrow     \nLeftrightarrow

Table 15: AMS Negated Arrows

                               \digamma     \varkappa

Table 16: AMS Greek

                            \beth     \daleth     \gimel

Table 17: AMS Hebrew

  h̄ \hbar              \hslash                 \vartriangle      \triangledown
      \square            \lozenge                \circledS     ^{}\angle
      \measuredangle     \nexists  ℧ \mho              \Finv^{u}
      \Game^{u}  \Bbbk^{u}  \backprime        \varnothing
      \blacktriangle     \blacktriangledown      \blacksquare      \blacklozenge
      \bigstar           \sphericalangle         \complement       \eth
      \diagup^{u}  \diagdown^{u}

u Not defined in amssymb.sty, define using the \newsymbol command. Table 18: AMS Miscellaneous

      \dotplus         \smallsetminus     \Cap                \Cup
      \barwedge        \veebar            \doublebarwedge     \boxminus
      \boxtimes        \boxdot            \boxplus            \divideontimes
      \ltimes          \rtimes            \leftthreetimes     \rightthreetimes
      \curlywedge      \curlyvee          \circleddash        \circledast
      \circledcirc     \centerdot         \intercal

Table 19: AMS Binary Operators

\leqq                   \leqslant               \eqslantless         \lesssim
\lessapprox             \approxeq               \lessdot             \lll
\lessgtr                \lesseqgtr              \lesseqqgtr          \doteqdot
\risingdotseq           \fallingdotseq          \backsim             \backsimeq
\subseteqq              \Subset  ⊏ \sqsubset            \preccurlyeq
\curlyeqprec            \precsim                \precapprox          \vartriangleleft
\trianglelefteq         \vDash                  \Vvdash              \smallsmile
\smallfrown             \bumpeq                 \Bumpeq              \geqq
\geqslant               \eqslantgtr             \gtrsim              \gtrapprox
\gtrdot                 \ggg                    \gtrless             \gtreqless
\gtreqqless             \eqcirc                 \circeq              \triangleq
\thicksim               \thickapprox            \supseteqq           \Supset
⊐ \sqsupset               \succcurlyeq            \curlyeqsucc         \succsim
\succapprox             \vartriangleright       \trianglerighteq     \Vdash
\shortmid               \shortparallel          \between             \pitchfork
\varpropto              \blacktriangleleft      \therefore           \backepsilon
\blacktriangleright     \because

Table 20: AMS Binary Relations

   \nless              \nleq                 \nleqslant         \nleqq
   \lneq               \lneqq                \lvertneqq         \lnsim
   \lnapprox           \nprec                \npreceq           \precnsim
   \precnapprox        \nsim                 \nshortmid         \nmid
   \nvdash             \nvDash               \ntriangleleft     \ntrianglelefteq
   \nsubseteq          \subsetneq            \varsubsetneq      \subsetneqq
   \varsubsetneqq      \ngtr                 \ngeq              \ngeqslant
   \ngeqq              \gneq                 \gneqq             \gvertneqq
   \gnsim              \gnapprox             \nsucc             \nsucceq
   \nsucceqq^{u}  \succnsim             \succnapprox       \ncong
   \nshortparallel     \nparallel            \nvDash            \nVDash
   \ntriangleright     \ntrianglerighteq     \nsupseteq         \nsupseteqq
   \supsetneq          \varsupsetneq         \supsetneqq        \varsupsetneqq

u Not defined in amssymb.sty, define using the \newsymbol command. Table 21: AMS Negated Binary Relations

                 newlfont/margid   oldlfont/nomargid Requried style option
       ABCdef \mathrm{ABCdef}   {\mathrm ABCdef}
       ABC  \cal{ABC}         {\cal ABC}
       𝒜ℬ𝒞  \EuScript{ABC}    {\EuScript ABCdef}  euscript
       ABCdef \frak{ABCdef}     {\frak ABCdef}      amsfonts or amssymb
       ABC  \Bbb{ABC}         {\Bbb ABC}          amsfonts or amssymb
       ABCdef \mathcyr{ABCdef}  {\mathcyr ABCdef}   cyrillic

Table 22: Math Alphabets